An isoseles triangle has a perimeter of 12cm.
One of its sides is 5cm.
What could the length of each of the other two sides be?

Answer by CharlesG2(834)

An isoseles triangle has a perimeter of 12cm.
One of its sides is 5cm.
What could the length of each of the other two sides be?
isosceles triangles have at least 2 of its 3 sides equal
side A + side B + side C = P perimeter = 12 cm
call sides A and B the equal sides
2A + C = 12
2 possibilities: either A and B are 5 cm which would make C = 2 cm
OR C = 5 cm which would make A and B equal to 3.5 cm
(2 * 3.5 = 7, and 7 + 5 = 12)